# Provably Safe Reinforcement Learning using Entropy Regularizer

### v2

@@ Abstract
We consider the problem of learning the optimal policy for Markov decision processes with safety constraints. We formulate the problem in a reach-avoid setup. Our goal is to design online reinforcement learning algorithms that ensure safety constraints with arbitrarily high probability during the learning phase. To this end, we first propose an algorithm based on the optimism in the face of uncertainty (OFU) principle. Based on the first algorithm, we propose our main algorithm, which utilizes entropy regularization. We investigate the finite-sample analysis of both of the algorithms and derive their regret bounds. We demonstrate that the inclusion of entropy regularization improves the regret and drastically controls the episode-to-episode variability that is inherent in OFU-based safe RL algorithms.
@@ 1 Introduction
Linear Programming Approach: CMDPs form the theoretical backbone of safe RL by extending standard MDPs to account for constraints on expected cumulative costs. The foundational work [4] introduces CMDPs using linear programming (LP) and Lagrangian duality. Building on the works presented in [4], [10] proposes a LP-based learning algorithms that achieve sublinear regret in online settings while respecting constraints. More recently, [22] develops algorithms that carefully balance exploration and constraint satisfaction. In [7], another algorithm is developed providing strong theoretical guarantees for safety and performance. However, all of the aforementioned works consider either a deterministic finite horizon or an infinite horizon.
Lagrangian and Primal-Dual Approach: Another key direction in Safe RL is the use of Lagrangian-based methods, which convert constrained optimization problems into unconstrained ones by introducing penalty terms. Reference [1] presents constrained policy optimization (CPO), a practical algorithm that adapts trust-region policy updates to enforce constraint satisfaction. Further, [23, 25, 24] introduce methods such as Triple-Q learning that combine improved convergence properties with enhanced empirical performance, particularly in episodic CMDPs.
@@
Our contributions: We consider the problem of learning an optimal policy for Markov decision processes (MDPs) subject to a safety constraint. Adopting a probabilistic notion of safety, namely $p$-safety [28], we formulate the problem within a reach–avoid framework. We assume that the state space of the MDP is partitioned into three disjoint subsets: a target or goal set comprising all goal states, an unsafe set comprising all unsafe states, and a living set consisting of all remaining states that are neither goal nor unsafe.
To formalize this setup, we employ a CMDP formulation [3]. We assume that the target set is terminal, meaning that the process terminates whenever the MDP reaches a target state, while the unsafe set and the living set are non-terminal. In contrast, our earlier work [15] assumed that the unsafe set was also terminal in order to reduce computational complexity.
@@ 2 Background and Problem Formulation
Let $p\in(0,1)$ and let $h$ be a scalar satisfying $0\leq h<p$. An action $a(x)\in\mathcal{A}$ is said to be $h$-safe if, when action $a(x)$ is applied at state $x$, the probability of transitioning to the unsafe set $U$ satisfies $\sum_{y\in U}P(x,a(x),y)\leq h.$
@@
[16] Let $U^{\prime}\subset H$. The set $U^{\prime}$ is called a proxy set if, for any policy $\pi\in\Delta(\mathcal{A})$ and any initial state $x_{0}\in H$, the stopping time $\tau_{U^{\prime}}$ is strictly smaller than $\tau_{U}$ almost surely, i.e., $\tau_{U^{\prime}}<\tau_{U}$ almost surely.
@@
Assumption 4 and Assumption 5 are required to have a finite stopping time and thus making the problem well-defined [29].
@@
For each state $x\in H\cup U$, there exist at least one safe action $a^{s}(x)$ with $h\leq\frac{p}{T_{max}}$.
Due to Assumption 5, the hitting time of set $E$ is finite almost surely. Further, set $E$ is accessible from set $H\cup U$ for any policy $\pi$. Hence, the stopping time $\tau_{E}$ is upper bounded by a positive real $T_{max}$ almost surely, i.e., $\mathds{P}^{x}_{\pi}[\tau_{E}\leq T_{max};P]=1$.
@@
In this section, we first develop an algorithm to learn the optimal policy for the learning problem (2). In our earlier work [15], we assumed that both the unsafe set $U$ and the target set $E$ are terminal. This assumption precludes transitions from the unsafe set $U$ to other states, thereby simplifying the underlying optimization problem. In this work, however, we adopt a more general and practically motivated modeling framework in which only the target set $E$ is terminal, while transitions from the unsafe set $U$ to the living set $H$ or directly to $E$ are permitted. This formulation captures a broad class of real-world systems in which safety violations are not necessarily absorbing events, but rather correspond to transient or recoverable conditions that incur increased risk. For instance, an autonomous robot may temporarily enter unsafe regions before corrective actions are applied, and clinical decision-making processes may involve adverse but reversible outcomes. By allowing non-absorbing unsafe states, the model more accurately reflects the underlying system dynamics and enables the development of learning algorithms that explicitly trade off performance and probabilistic safety.
@@
$\pi^{*}(a|x)=\frac{{\gamma}^{*}(x,a)+\beta^{*}(x,a)}{\sum_{a\in\mathcal{A}}\left({\gamma}^{*}(x,a)+\beta^{*}(x,a)\right)}.$ (5)
@@
1:Input: Safety parameter $p\in(0,1)$, initial state $x_{0}$, confidence parameter $\delta\in(0,1)$, one $h$-safe action $a^{safe}(x^{\prime})$ for each proxy state, cost $c(x,a)$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$
@@
Considering all the possible realizations, we obtain equality $(a)$ as follows:
@@
$\begin{split}&\overset{}{\leq}T_{max}\Big(q\ h+\frac{(|\mathcal{A}|-1)}{(|\mathcal{A}|-1)}(1-q)\Big)\\ &=T_{max}\Big(1-q\ (1-h))\Big)\\ &\overset{(e)}{\leq}p.\end{split}$
Equality $(e)$ follows from the definition of the safe action. The following result shows that the policy $\pi_{k}$ computed by the PSRL algorithm is safe with arbitrarily high probability. The proof follows similar arguments to those used in the proof of Theorem $2$ in [15].
@@
$R(K):=\sum_{k=1}^{K}\left(V^{P}_{\pi_{k}}(c,x_{0})-V^{P}_{\pi^{*}}(c,x_{0})\right).$
Moreover, we also define a per-episode objective regret and per-episode constraint regret defined as:
$\begin{split}&\mathcal{R}_{k}:=V^{P}_{\pi_{k}}(c,x_{0})-V^{P}_{\pi^{*}}(c,x_{0}),\\ &\mathcal{C}_{k}:=S^{P}_{\pi_{k}}(x_{0})-S^{P}_{\pi^{*}}(x_{0})\end{split}$
@@
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the cumulative objective regret of the PSRL algorithm after $K$ episodes satisfies
@@ 4 Entropy regularized $p$-Safe RL (ER-PSRL) Algorithm
In this section, we present our main algorithm, entropy-regularized $p$-safe reinforcement learning (ER–PSRL), which enhances the optimization problem underlying the PSRL algorithm by incorporating an entropy regularization term.
@@
  • With a linear objective, the LP optimum lies at an extreme point. Small perturbations in $\hat{P}_{k}$ can switch the active set, moving the solution to a different vertex. Hence, the solution $(\hat{h}_{k},\hat{g}_{k})$ are non-Lipschitz with respect to $(\hat{P}_{k},\tilde{c}_{k})$, which leads to unstable policy updates and large episode-to-episode fluctuations in the per-episode objective regret, even when the underlying estimation noise is small.
@@
We define the entropy function $H$ with the following insight. From the definition of the occupation measure $\gamma(x,a)$ and $\beta(x,a)$ it is apparent that $\tilde{\xi}(x,a)\leq T_{max}$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$. We consider $\tilde{\xi}(x,a)/T_{max}$ instead of $\tilde{\xi}(x,a)$ as this will make $H\geq 0$ for all possible solutions $\tilde{\xi}$. By introducing the parameter $\eta>0$ in the entropy function $H$ helps us in deriving the desired regret bound later.
The pseudo-code for the ER-PSRL algorithm is as given in Algorithm $2$.
@@
1:Input: Safety parameter $p\in(0,1)$, initial state $x_{0}$, confidence parameter $\delta\in(0,1)$, one safe action $a^{safe}(x^{\prime})$ for each proxy state, cost $c(x,a)$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$
@@
The set of feasible solutions of the optimization is exactly the same as the extended LP (11)-(12). Thus, since the policy designed using the PSRL algorithm is safe with an arbitrary high probability, so is the policy designed using the ER-PSRL algorithm. We formally present this result as follows.
Theorem 12 (The ER–PSRL algorithm is $p$-safe).
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the policy $\pi_{k}$ computed by the ER–PSRL algorithm is safe, i.e., for the initial state $x_{0}$, the associated safety function satisfies
@@
The regret bound for ER-PSRL algorithm is as given below.
@@
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the cumulative objective regret of the ER-PSRL algorithm after $K$ episodes satisfies
@@
The ER-PSRL algorithm removes the explicit dependence of the regret bound on $(p-p^{s})^{-1}$. Since $0<p-p^{s}<1$, the regret bound with the OFU-based PSRL is always no smaller than the entropy-regularized bound. In particular, when the baseline policy operates close to the safety threshold ($p^{s}\approx p$), the regret bound of Algorithm 1 deteriorates significantly, whereas ER-PSRL maintains a stable $\tilde{\mathcal{O}}\!\left(|\mathcal{X}|\sqrt{|\mathcal{A}|T_{\max}^{3}K}\right)$ rate. This highlights the role of entropy regularization in mitigating the adverse effect of a small safety margin while preserving sublinear regret.
@@ 5 Simulation results
Figure 4: Cumulative objective regret for ER-PSRL
@@ 6 Conclusion and Future Works
In this paper, we proposed two reinforcement learning algorithms that guarantee probabilistic safety with arbitrarily high confidence. We provided a finite-sample analysis of both algorithms and established explicit regret bounds. Our analysis revealed that the LP-based PSRL algorithm, while theoretically sound, suffers from significant episode-to-episode variability and insufficient exploration, which can degrade its empirical performance. To address these limitations, we introduced the entropy-regularized PSRL (ER-PSRL) algorithm which reduces high inter-episode variability and promotes systematic exploration in a principled manner. We theoretically demonstrated that ER-PSRL achieves a strictly improved regret bound compared to the LP-based PSRL algorithm.
@@
For the rest of the proofs and results are based on a stopped Markov decision process defined from the original MDP.
@@
Suppose $X^{\tau}_{k,t}$ and $A^{\tau}_{k,t}$ represent the state and the action at time $t$ in episode $k$. For episode $k$, we define the following modified cost:
@@
For all time steps $t\geq\tau_{E}$ and all episodes $k$, the policy $\pi$ is defined as a deterministic policy, i.e., $\pi(X^{\tau}_{k,t})=a^{\prime}$ where $a^{\prime}$ is a randomly fixed action. We also define $\sigma_{k,t}(x,a)$, the occupation measure for stopped MDP ${M}^{\tau}$ as follows:
@@
The occupation measure $\sigma_{k,t}(x,a)$ satisfies the following property:
@@
For stopped MDP ${M}^{\tau}$, we use $\mathcal{F}^{\tau}_{k}$ to denote the filtration up to the the episode $k$ consisting of all state-action pairs visited until the end of episode $k$, i.e., $\mathcal{F}^{\tau}_{k}:\{X^{\tau}_{k^{\prime},t},A^{\tau}_{k^{\prime},t},k^{\prime}\in[1,k]\ \text{and}\ t\leq\tau_{E}\}$. We also define the following variables for each episode $k$:
@@
It is easy to see that ${N}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t})=N_{k}(X_{k,t},A_{k,t})$ for $t\leq\tau_{E}-1$. Since, target set $E$ is terminal so from the construction of the stopped MDP $M^{\tau}$, ${N}^{\tau}_{k}(x,a)=N_{k}(x,a)$ for all $k\in[1,K]$, $a\in\mathcal{A}$ and $x\in H\cup U$. We define the optimistic transition probability $\tilde{P}^{\tau}_{k}$ for the stopped MDP $M^{\tau}$ as follows:
$\tilde{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},X^{\tau}_{k,t+1}):=\begin{cases}&\tilde{P}_{k}(X_{k,t},A_{k,t},X_{k,t+1});\ \forall t\leq\tau_{E}-1\\ &P^{\tau}(X_{k,t},A_{k,t},X_{k,t+1});\ \forall t>\tau_{E}-1\end{cases}$
@@
With probability at least $(1-2\delta)$, the following is true for all $k\in\{1,2,...,K\}$:
@@
    For all $(x,a,y)\in\mathcal{X}\times\mathcal{A}\times\mathcal{\mathcal{X}}$,
     $\begin{split}\text{ }|P^{\tau}(x,a,y)-\hat{P}^{\tau}_{k}(x,a,y)|\leq\epsilon^{\tau}_{k}(x,a,y),\end{split}$
@@
    For all $(x,a,y)\in\mathcal{X}\times\mathcal{A}\times\mathcal{\mathcal{X}}$,
    $\begin{split}|P^{\tau}(x,a,y)-\tilde{P}^{\tau}_{k}(x,a,y)|\leq 2\epsilon^{\tau}_{k}(x,a,y).\end{split}$
@@
    For all $(x,a)\in(H\cup U)\times\mathcal{A}$,
    $\kappa(x,a)\leq\hat{\kappa}_{k}(x,a)+\hat{\epsilon}^{\tau}_{k}(x,a),$
@@
For all $(x,a,y)\in E\times\mathcal{A}\times\mathcal{X}$, claim $(i)$ and $(ii)$ are true from the definition. For all $(x,a,y)\in(H\cup U)\times\mathcal{A}\times\mathcal{X}$, claim $(i)$, $(ii)$ and $(iii)$ are proved in [15].
@@
Suppose OFU (7) is feasible. Then, the learned policy ${\pi}_{k}$ is an optimistic policy for the original MDP $M$, meaning that ${V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$. Further for the stopped MDP $M^{\tau}$, ${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P^{\tau}}_{{\pi}^{*}}({c},x_{0})$.
@@
$\tilde{S}_{\check{\pi}}^{\check{P}}(x_{0}):=\mathds{E}_{\check{\pi}}^{x,\check{P}}\sum_{t=0}^{\tau-1}\big(\tilde{\kappa}(X_{t},A_{t})+3\hat{\epsilon}_{k}(X_{t},A_{t})\big),$

### v3

# Provably Safe Reinforcement Learning for Stochastic Reach-Avoid Problems with Entropy Regularization

@@ Abstract
We consider the problem of learning the optimal policy for Markov decision processes with safety constraints. We formulate the problem in a stochastic reach-avoid setup. Our goal is to design online reinforcement learning algorithms that ensure safety constraints with arbitrarily high probability during the learning phase. To this end, we first propose an algorithm based on the optimism in the face of uncertainty (OFU) principle. Based on the first algorithm, we propose our main algorithm, which utilizes entropy regularization. We investigate the finite-sample analysis of both algorithms and derive their regret bounds. We demonstrate that the inclusion of entropy regularization improves the regret and significantly controls the episode-to-episode variability inherent in OFU-based safe RL algorithms.
@@ 1 Introduction
Linear Programming Approach: CMDPs form the theoretical backbone of safe RL by extending standard MDPs to account for constraints on expected cumulative costs. The foundational work [4] introduces CMDPs using linear programming (LP) and Lagrangian duality. Building on the works presented in [4], [10] proposes a LP-based learning algorithms that achieve sublinear regret in online settings while respecting constraints. More recently, [22] develops algorithms that carefully balance exploration and constraint satisfaction. In [7], another algorithm is developed that ensures safety throughout the learning phase and provides a better theoretical regret bound. However, all of the aforementioned works consider either a deterministic finite horizon or an infinite horizon.
Lagrangian and Primal-Dual Approach: Another key direction in safe RL is the use of Lagrangian-based methods, which convert constrained optimization problems into unconstrained ones by introducing penalty terms. Reference [1] presents constrained policy optimization (CPO), a practical algorithm that adapts trust-region policy updates to enforce constraint satisfaction. Further, [24, 23] introduce Q-learning-based methods that combine improved convergence properties with enhanced empirical performance, particularly in episodic CMDPs.
@@
Our contributions: We consider the problem of learning an optimal policy for Markov decision processes (MDPs) subject to a safety constraint. Adopting a probabilistic notion of safety, namely $p$-safety [27], we formulate the problem within a stochastic reach–avoid framework. We assume that the state space of the MDP is partitioned into three disjoint subsets: a target or goal set comprising all goal states, an unsafe set comprising all unsafe states, and a living set consisting of all remaining states that are neither goal nor unsafe.
To formalize this setup, we employ a CMDP formulation [3]. We assume that the target set is terminal, meaning that the process terminates whenever the MDP reaches a target state, while the unsafe set and the living set are non-terminal. In contrast, our earlier work [15] assumes that the unsafe set is also terminal to reduce computational complexity. The setup considered in [15] precludes transitions from the unsafe set to other states, thereby simplifying the underlying optimization problem. In this work, however, we adopt a more general and practically motivated modeling framework in which only the target set is terminal, while transitions from the unsafe set to the living set or directly to the goal set are permitted. This formulation captures a broad class of real-world systems in which safety violations are not necessarily absorbing events, but rather correspond to transient or recoverable conditions that incur increased risk. For instance, an autonomous robot may temporarily enter unsafe regions before corrective actions are applied, and clinical decision-making processes may involve adverse but reversible outcomes. By allowing non-absorbing unsafe states, the model more accurately reflects the underlying system dynamics and enables the development of learning algorithms that explicitly trade off performance and probabilistic safety.
@@ 2 Background and Problem Formulation
Let $p\in(0,1)$ and let $h$ be a scalar satisfying $0\leq h<p$. An action $a(x)\in\mathcal{A}$ is said to be safe if, when action $a(x)$ is applied at state $x$, the probability of transitioning to the unsafe set $U$ satisfies $\sum_{y\in U}P(x,a(x),y)\leq h.$
@@
[16] Let $U^{\prime}\subset H$. The set $U^{\prime}$ is said to be a proxy set if the following conditions hold:
  • For any policy $\pi\in\Delta(\mathcal{A})$ and any initial state $x_{0}\in H$, the stopping time $\tau_{U^{\prime}}$ occurs strictly before $\tau_{U}$ almost surely, i.e., $\tau_{U^{\prime}}<\tau_{U}$ almost surely.
  • From any state $x\in U^{\prime}$, the unsafe set $U$ is reached in a single transition with probability one, i.e.:
    $\sum_{a\in\mathcal{A}}\sum_{y\in U}P(x,a,y)=1.$
@@
Assumption 4 and Assumption 5 are required to have a finite stopping time for all policies (deterministic or stochastic), thus making the problem well-defined [28].
Due to Assumption 5, the hitting time of set $E$ is finite almost surely. Further, set $E$ is accessible from set $H\cup U$ for any policy $\pi$. Hence, the stopping time $\tau_{E}$ is upper bounded by a positive real $T_{max}$ almost surely, i.e., $\mathds{P}^{x_{0}}_{\pi}[\tau_{E}\leq T_{max};P]=1$.
@@
The following assumption guarantees the existence of a safe baseline policy for the proposed algorithms.
@@
For every state $x\in U^{\prime}$, there exists at least one safe action $a^{s}(x)$ satisfying $h\leq\frac{p}{T_{\max}}$.
@@
In this section, we first develop an algorithm to learn the optimal policy for the learning problem (2).
@@
$\pi^{*}(a|x)=\frac{{\gamma}^{*}(x,a)+\beta^{*}(x,a)}{\sum_{a\in\mathcal{A}}\left({\gamma}^{*}(x,a)+\beta^{*}(x,a)\right)},$ (5)
@@
1:Input: Safety parameter $p\in(0,1)$, initial state $x_{0}$, confidence parameter $\delta\in(0,1)$, one safe action $a^{s}(x^{\prime})$ for each proxy state, cost $c(x,a)$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$
@@
Considering all the possible realizations, we obtain inequality $(a)$ as follows:
@@
$\begin{split}&\overset{}{\leq}T_{max}\Big(q\ h+\frac{(|\mathcal{A}|-1)}{(|\mathcal{A}|-1)}(1-q)\Big)\\ &=T_{max}\Big(1-q\ (1-h))\Big)\\ &\overset{(a)}{\leq}p.\end{split}$
Inequality $(a)$ follows from the definition of the safe action. ∎
The following result establishes the safety guarantee of the pSRL algorithm. The proof proceeds along similar lines to the arguments used in the proof of Theorem 2 in [15].
@@
We define the per-episode objective regret and per-episode constraint regret as:
$\begin{split}&\mathcal{R}_{k}:=V^{P}_{\pi_{k}}(c,x_{0})-V^{P}_{\pi^{*}}(c,x_{0}),\\ &\mathcal{C}_{k}:=S^{P}_{\pi_{k}}(x_{0})-S^{P}_{\pi^{*}}(x_{0}).\end{split}$
@@
$R(K):=\sum_{k=1}^{K}\mathcal{R}_{k}.$
@@
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-4\delta)$, the cumulative objective regret of the pSRL algorithm after $K$ episodes satisfies
@@ 4 Entropy regularized $p$-Safe RL (ER-pSRL) Algorithm
In this section, we present our main algorithm, entropy-regularized $p$-safe reinforcement learning (ER-pSRL), which enhances the optimization problem underlying the pSRL algorithm by incorporating an entropy regularization term.
@@
  • With a linear objective, the optimum of an LP is attained at an extreme point of the feasible region. Consequently, even small perturbations in $\hat{P}_{k}$ can change the active constraint set and move the optimizer to a different vertex. As a result, the solution $(\hat{h}_{k},\hat{g}_{k})$ may vary significantly in response to minor changes in $(\hat{P}_{k},\tilde{c}_{k})$, leading to unstable policy updates and large episode-to-episode fluctuations in the per-episode objective regret, even when the underlying estimation noise is small.
@@
We introduce the entropy function $H$ based on the following considerations. From the definition of the occupation measures $\gamma(x,a)$ and $\beta(x,a)$, it follows that $\tilde{\xi}(x,a)\leq T_{\max}$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$. Accordingly, we normalize the occupation measure by $T_{\max}$ and define the entropy in terms of $\tilde{\xi}(x,a)/T_{\max}$ rather than $\tilde{\xi}(x,a)$, which ensures that $H\geq 0$ for all feasible solutions $\tilde{\xi}$. Furthermore, we introduce a small parameter $\eta\in(0,1)$, chosen to be sufficiently smaller than typical values of $\tilde{\xi}$. While the inclusion of $\eta$ has a negligible effect on the entropy term itself, it plays a crucial technical role in the derivation of the desired regret bound.
The pseudo-code for the ER-pSRL algorithm is given in Algorithm $2$.
@@
1:Input: Safety parameter $p\in(0,1)$, initial state $x_{0}$, confidence parameter $\delta\in(0,1)$, one safe action $a^{s}(x^{\prime})$ for each proxy state, cost $c(x,a)$ for all $(x,a)\in\mathcal{X}\times\mathcal{A}$
@@
The set of feasible solutions of the optimization is exactly the same as the extended LP (11)-(12). Thus, since the policy designed using the pSRL algorithm is safe with an arbitrarily high probability, so is the policy designed using the ER-pSRL algorithm. We formally present this result as follows.
Theorem 12 (The ER-pSRL algorithm is $p$-safe).
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-2\delta)$, the policy $\pi_{k}$ computed by the ER-pSRL algorithm is safe, i.e., for the initial state $x_{0}$, the associated safety function satisfies
@@
The regret bound for the ER-pSRL algorithm is as given below.
@@
Let $\delta\in\bigl(0,\tfrac{1}{2}\bigr)$. Then, with probability at least $(1-4\delta)$, the cumulative objective regret of the ER-pSRL algorithm after $K$ episodes satisfies
@@
The ER-pSRL algorithm eliminates the explicit dependence of the regret bound on $(p-p^{s})^{-1}$. Since $0<p-p^{s}<1$, the regret bound associated with the OFU-based pSRL algorithm is always no smaller than its entropy-regularized counterpart. In particular, when the safe baseline policy operates close to the safety threshold ($p^{s}\approx p$), the regret bound for pSRL deteriorates sharply and becomes vacuous as $p\to p^{s}$. In contrast, ER-pSRL preserves a stable regret bound of order $\tilde{\mathcal{O}}\!\left(|\mathcal{X}|\sqrt{|\mathcal{A}|T_{\max}^{3}K}\right)$, thereby demonstrating the effectiveness of entropy regularization in mitigating the adverse effects of a vanishing safety margin while maintaining sublinear regret.
@@ 5 Simulation results
Figure 4: Cumulative objective regret for the ER-pSRL algorithm
@@ 6 Conclusion and Future Works
In this paper, we proposed two reinforcement learning algorithms that guarantee probabilistic safety with arbitrarily high confidence. We provided a finite-sample analysis of both algorithms and established explicit regret bounds. Our analysis revealed that the LP-based pSRL algorithm, while theoretically sound, suffers from significant episode-to-episode variability and insufficient exploration, which can degrade its empirical performance. To address these limitations, we introduced the ER-pSRL algorithm, which reduces high inter-episode variability and promotes systematic exploration in a principled manner. We theoretically demonstrated that ER-pSRL achieves a strictly improved regret bound compared to the LP-based pSRL algorithm.
@@
The rest of the proofs and results are based on a stopped Markov decision process defined from the original MDP.
@@
Suppose $X^{\tau}_{k,t}$ and $A^{\tau}_{k,t}$ represent the state and the action seen by the algorithm at time $t$ in episode $k$. For episode $k$, we define the following modified cost:
@@
For all time steps $t\geq\tau_{E}$ and all episodes $k$, the policy $\pi$ is defined as a deterministic policy, i.e., $\pi(X^{\tau}_{k,t})=a^{\prime}$ where $a^{\prime}$ is a any fixed action. We also define $\sigma_{k,t}(x,a)$, the occupation measure for stopped MDP ${M}^{\tau}$ as follows:
@@
The occupation measure $\sigma_{k,t}(x,a)$ satisfies the following property for all $t\geq 1$:
@@
For stopped MDP ${M}^{\tau}$, we use $\mathcal{F}^{\tau}_{k}$ to denote the filtration up to the the episode $k$ consisting of all state-action pairs visited until the end of episode $k$, i.e., $\mathcal{F}^{\tau}_{k}:\{X^{\tau}_{k^{\prime},t},A^{\tau}_{k^{\prime},t},k^{\prime}\in[1,k]\ \text{and}\ t\leq T_{max}-1\}$. We also define the following variables for each episode $k$:
@@
Since, target set $E$ is terminal so from the construction of the stopped MDP $M^{\tau}$, ${N}^{\tau}_{k}(x,a)=N_{k}(x,a)$ for all $k\in[1,K]$, $a\in\mathcal{A}$ and $x\in H\cup U$. We define the optimistic transition probability $\tilde{P}^{\tau}_{k}$ for the stopped MDP $M^{\tau}$ as follows:
$\tilde{P}^{\tau}_{k}(X^{\tau}_{k,t},A^{\tau}_{k,t},X^{\tau}_{k,t+1}):=\begin{cases}&\tilde{P}_{k}(X_{k,t},A_{k,t},X_{k,t+1});\ \forall t\leq\tau_{E}-1\\ &P^{\tau}(X_{k,t},A_{k,t},X_{k,t+1});\ \forall t>\tau_{E}-1\end{cases}$ (20)
@@
With probability at least $(1-2\delta)$, the following is true for all $k\in\{1,2,...,K\}$ and for all $(x,a,y)\in\mathcal{X}\times\mathcal{A}\times\mathcal{\mathcal{X}}$:
@@
    $|P^{\tau}(x,a,y)-\hat{P}^{\tau}_{k}(x,a,y)|\leq\epsilon^{\tau}_{k}(x,a,y)$.
@@
    $|P^{\tau}(x,a,y)-\tilde{P}^{\tau}_{k}(x,a,y)|\leq 2\epsilon^{\tau}_{k}(x,a,y).$
@@
    $|\kappa(x,a)-\hat{\kappa}_{k}(x,a)|\leq\hat{\epsilon}^{\tau}_{k}(x,a),$
@@
For all $(x,a,y)\in E\times\mathcal{A}\times\mathcal{X}$, claim $(i)$ and $(ii)$ are true from the definition. For all $(x,a,y)\in(H\cup U)\times\mathcal{A}\times\mathcal{X}$, claim $(i)$, $(ii)$ and $(iii)$ are proved in [15]. ∎
@@
Suppose OFU (7)-(8) is feasible. Then, the learned policy ${\pi}_{k}$ is an optimistic policy for the original MDP $M$, meaning that ${V}^{\tilde{P}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P}_{{\pi}^{*}}({c},x_{0})$. Further for the stopped MDP $M^{\tau}$, ${V}^{\tilde{P}^{\tau}_{k}}_{{\pi}_{k}}(\tilde{c}_{k},x_{0})\leq{V}^{P^{\tau}}_{{\pi}^{*}}({c},x_{0})$.
@@
$\tilde{S}_{\check{\pi}}^{\check{P}}(x_{0}):=\mathds{E}_{\check{\pi}}^{x,\check{P}}\sum_{t=0}^{\tau-1}\big(\hat{\kappa}(X_{t},A_{t})+3\hat{\epsilon}_{k}(X_{t},A_{t})\big),$